### v2

# Progressive Quantum Algorithm for Maximum Independent Set with Quantum Alternating Operator Ansatz

Abstract
Recently, Hadfield et al. proposed the Quantum Alternating Operator Ansatz (QAOA+) to tackle Constrained Combinatorial Optimization Problems (CCOPs). This paper proposes a Progressive Quantum Algorithm (PQA) to reduce the required qubits in solving the Maximum Independent Set (MIS) problem using QAOA+ ansatz. PQA constructs a subgraph that contains the MIS solution of the target graph $G$ and then solves the MIS problem on this subgraph to obtain the solution for $G$. To induce such a subgraph, PQA starts with a small-scale initial subgraph and progressively expands its graph size utilizing designed expansion rules. After each expansion, PQA solves the MIS problem on the current subgraph using QAOA+. In each run, PQA repeats the graph expansion and solving process until a predefined stopping condition is reached. Simulation results demonstrate that to achieve an optimal approximation ratio of 0.95, PQA requires only $5.5565\%$ ($15.4\%$) of the qubits and $10.695\%$ ($7.23\%$) of the runtime compared with QAOA+ on Erdős-Rényi (regular) graphs, highlighting the efficiency of PQA.
pacs:
Valid PACS appear here
preprint: APS/123-QED
I INTRODUCTION
Over recent years, quantum computing has emerged as a promising computational paradigm, demonstrating its potential for addressing complex problems such as database searching Grover (1997), prime factorization Shor (1999), data dimension reduction Lloyd et al. (2014); Yu et al. (2019); Pan et al. (2020); Wan et al. (2018) and solving equations Harrow et al. (2009). However, the practical implementation of these quantum algorithms is hindered by noise and the limited qubit capacity of existing Noisy Intermediate-Scale Quantum (NISQ) devices Preskill (2018); Cerezo et al. (2021); Bharti et al. (2022). Therefore, hybrid quantum-classical algorithms Grimsley et al. (2019); Tang et al. (2021); Zhu et al. (2019); Du et al. (2022); He et al. (2024); Song et al. (2024); Wu et al. (2023a, b); Wang et al. (2021); Liu et al. (2021) have been proposed to harness the capabilities of NISQ devices more effectively. The Quantum Approximate Optimization Algorithm (QAOA) Farhi et al. (2014); Bravyi et al. (2020); Zhou et al. (2020); Sack and Serbyn (2021); Streif and Leib (2020); Zhu et al. (2022); Ni et al. (2024); Zhou et al. (2023); Wurtz and Lykov (2021) is a hybrid quantum-classical algorithm with broad applications. Specifically, it has been used to solve the exact cover Vikstål et al. (2020), vertex cover Zhang et al. (2022), and other Combinatorial Optimization Problems (COPs) Wurtz and Love (2021); Ruan et al. (2023); Brandhofer et al. (2022); Tomesh et al. (2023); Finžgar et al. (2023). Constrained COPs (CCOPs) are a crucial subclass of combinatorial optimization, and solving CCOPs requires not only finding the optimal solution but also satisfying specific constraints, making them more complex and challenging compared with Unconstrained COPs (UCOPs) Karp (2010). To address these challenges for solving CCOPs, Hadfield et al. Hadfield et al. (2019) proposed an improved version of the original QAOA Farhi et al. (2014), known as QAOA+. QAOA+ constructs a Hilbert space that only contains feasible solutions to the problem, thereby reducing the solution search space compared with the original QAOA Ruan et al. (2023). Currently, QAOA+ has been applied to solve Maximum Independent Set (MIS) Saleem (2020), minimum exact cover Wang et al. (2023) and other CCOPs Hadfield et al. (2017); Wang et al. (2020); He et al. (2023). Given the limited resources of NISQ devices, developing heuristic strategies to reduce the qubit requirements of QAOA+ in solving CCOPs becomes essential. Recently, Tomesh et al. Tomesh et al. (2022a) proposed Quantum Local Search (QLS) with the QAOA+ ansatz and studied its performance on the MIS problem. Different from directly solving the MIS problem using QAOA+ (DS-QAOA+) Saleem (2020), QLS constructs a global solution by iterating through many local subproblems. Specifically, QLS constructs multiple local subgraphs (that may be of indeterminate graph size) based on certain strategies and solves the MIS problem on these subgraphs using QAOA+ to obtain local solutions. These local solutions are combined to form a global solution using classical methods. For DS-QAOA+, several qubits equal to the size of the graph are required. By breaking the problem into smaller subgraphs, QLS reduces the number of qubits needed for the MIS problem. Inspired by the Proactively Incremental Learning (PIL) algorithm Li et al. (2023), a new method called the Progressive Quantum Algorithm (PQA) is proposed to reduce the required qubits for solving the MIS problem. PQA aims to construct a subgraph containing the MIS solution of the original problem and then solve the MIS problem on this subgraph to obtain the final solution. Specifically, PQA starts with a small, sparse initial subgraph and progressively expands its size following rules distinct from QLS. After each expansion, PQA solves the MIS problem on the newly generated subgraph using QAOA+. In each run, PQA repeats the graph expansion and solving process until a predefined stopping condition is reached. By doing so, PQA may solve the original MIS problem on a smaller-scale subgraph compared with DS-QAOA+, and it does not require reconstructing the global solution from multiple local solutions compared with QLS. In our simulations, we compare the total qubits and runtime consumed by PQA, QLS, and DS-QAOA+ on multiple Erdős-Rényi (ER) and $k(=2,3)$-regular graphs when they achieve the same optimal approximation ratio (OAR). Across all tested graphs, PQA usually achieves the same OAR with fewer qubits and runtime than DS-QAOA+ and QLS. Specifically, for an OAR of 0.95, PQA uses $5.5565\%$ ($26.77\%$) of the qubits required by DS-QAOA+ (QLS) on ER graphs, and $15.4\%$ ($52.605\%$) on regular graphs, respectively. Additionally, PQA reaches the same OAR in $10.695\%$ ($75.365\%$) of the runtime required by DS-QAOA+ (QLS) on ER graphs, and $7.23\%$ ($63.91\%$) on regular graphs. The MIS problem is one of the most widely studied CCOPs, partly due to its broad applicability in various domains and its equivalence to the minimum vertex cover and maximum clique. The efficiency of PQA in solving the MIS problem not only may enhance the practicality of quantum algorithms but also pave the way for solving larger-scale problems in the future. This paper is structured as follows: Section II serves as the foundational segment, providing background knowledge on the MIS problem, QAOA, and its extension, QAOA+. Following this, in Section III, we delineate the progressive quantum algorithm, emphasizing the construction rules of the desired derived subgraph. Subsequently, in Section IV, we present a comprehensive set of numerical simulation experiments aimed at showcasing the quantum resource efficiency of PQA over DS-QAOA+ and QLS. Finally, in Section V, we offer a succinct conclusion and discussion to encapsulate the key findings of our study and the next research direction.
II PRELIMINARIES
In this section, we introduce some preliminaries to help readers understand our work better.
II.1 Review of MIS
MIS is defined on a graph $G=(V,E)$, where $V=\left\{{0,...,n-1}\right\}$ represents the set of vertices, and $E=\left\{\left(u,v\right)\right\}$ represents the set of edges, with the two vertices $u$ and $v$ on the edge $(u,v)$ being adjacent. In this paper, we denote the number of vertices and edges as $n$ and $m$, respectively. A vertex subset $V_{s}$ is called an independent set of $G$ if no two vertices in $V_{s}$ are adjacent. The MIS is the vertex subset with the largest number of vertices among all independent sets, and the number of vertices in the MIS is known as the independence number, denoted as $\beta(G)$ Karp (2010). While the solution to MIS for a graph $G$ may not be unique, the independence number is always determined. For any target graph $G$, the independent vertex subsets in an induced subgraph $G_{q}$ of $G$ are feasible solutions for $G$. There is at least one induced subgraph $G_{q}=(V_{q},E_{q})$ that has an identical MIS solution to $G$, where $V_{q}\subseteq V$ and $E_{q}\subseteq E$. As illustrated in FIG. 1, the right graph is the induced subgraph of the left graph, and they have the same MIS solution $\left\{1,2,3\right\}$, but the MIS solution $\left\{1,2,4\right\}$ to FIG 1 is not the MIS solution of FIG. 1 because the node $4$ is not in this induced subgraph. This implies that the MIS solution of the target graph may not always be a solution for its induced subgraph, and an independent set in $G$ may not be feasible for $G_{q}$.
Figure 1: Examples of MIS. (a) A target graph $G$ whose feasible solutions are $\varnothing$, $\left\{0\right\}$, $\left\{1\right\}$, $\left\{2\right\}$, $\left\{3\right\}$, $\left\{4\right\}$, $\left\{1,2\right\}$, $\left\{1,3\right\}$, $\left\{1,4\right\}$, $\left\{2,3\right\}$, $\left\{2,4\right\}$, $\left\{1,2,3\right\}$ and $\left\{1,2,4\right\}$. The solutions of MIS are $\left\{1,2,3\right\}$ and $\left\{1,2,4\right\}$, where the MIS solutions of $G$ are not unique, but the independence number identically equals $3$. (b) An induced subgraph of $G$ and its MIS solution is $\left\{1,2,3\right\}$. The feasible solutions of $G_{q}$ are $\varnothing$, $\left\{0\right\}$, $\left\{1\right\}$, $\left\{2\right\}$, $\left\{3\right\}$, $\left\{1,2\right\}$, $\left\{1,3\right\}$, $\left\{2,3\right\}$, $\left\{1,2,3\right\}$. We respectively denote the feasible space including all feasible solutions of $G_{q}$ and $G$ as $S(G_{q})$ and $S(G)$, where $S(G_{q})\subseteq S(G)$.
In this paper, we utilize $x_{u}=1$ to represent that the vertex $u$ is in the vertex subset, otherwise, $x_{u}=0$, where $u=0,1,...,n-1$. Each way of vertex partition corresponds to a unique bit string $x$, and there are $2^{n}$ forms for $x$ in total, where $x=x_{0}x_{1}\dots x_{n-1}$. The goal for MIS is to maximize the objective function
$C_{1}(x)=\sum_{u=0}^{n-1}x_{u}$ (1)
when the independence constraint $\sum_{(u,v)\in E}x_{u}x_{v}=0$ is met. This constraint can be incorporated into the objective function using the Lagrange multiplier method Brady and Hadfield (2023), thus converting the constrained problem to unconstrained problem. The obtained classical cost function is
$\max C(x)=\sum_{u=0}^{n-1}x_{u}-\lambda\sum_{(u,v)\in E}x_{u}x_{v},$ (2)
where the value of $\lambda$ is determined through additional hyper-parameter optimization steps Finžgar et al. (2023), and $\lambda>1$ can impose penalties on those terms that violate independence constraints Lucas (2014), making the maximal cost function value correspond to the exact MIS solution instead of other feasible or infeasible bit strings. Here, $C_{\text{max}}$ equals the independence number $\beta(G)$.
II.2 Review of QAOA
Inspired by quantum adiabatic evolution Albash and Lidar (2018), QAOA encodes the solution of the problem into the ground state of the target Hamiltonian $H_{C}$ and aims to gradually evolve towards this ground state through a parameterized quantum circuit (PQC). This evolution starts from the ground state $|s\rangle$ of the initial Hamiltonian $H_{B}$, with a quantum circuit built using $p$-level QAOA ansatz. Notably, the initial state $|s\rangle$ is required to be trivial to implement. That is, $|s\rangle$ can be prepared by a constant-depth quantum circuit. For the original QAOA, one level QAOA ansatz consists of two variational QAOA parameters (denoted as $\gamma_{i}$ and $\beta_{i}$) and two unitaries $e^{-i\gamma_{i}H_{C}}$ and $e^{-i\beta_{i}H_{B}}$, where $i=1,2,\cdots,p$, and $p$ is the level depth. Generally, the initial Hamiltonian is chosen as $H_{B}=-\sum_{u=0}^{n-1}\sigma_{u}^{x}$ because its ground state $|s\rangle=|+\rangle^{\otimes n}$ can be trivial to prepare. Here, $\sigma_{u}^{x}$ refers to the Pauli-X operator applied to the $u-$th qubit, effectively flipping its state. To encode the MIS solution into the ground state of $H_{C}$, the cost function $-C(x)$ is transformed into the target Hamiltonian
$H_{C}=\sum_{u=0}^{n-1}\frac{\sigma_{u}^{z}-I}{2}+\lambda\sum_{(u,v)\in E}\frac{I-\sigma_{u}^{z}-\sigma_{v}^{z}+\sigma_{u}^{z}\sigma_{v}^{z}}{4}$ (3)
by transforming each binary variable $x_{u}$ to a quantum spin $\frac{I-\sigma_{u}^{z}}{2}$ Farhi et al. (2014), where $\sigma_{u}^{z}$ refers that applying Pauli-Z to the $u$-th qubit. The ground state of $H_{C}$ can be approximately obtained by alternately applying unitaries operators $e^{-i\gamma_{i}H_{C}}$ and $e^{-i\beta_{i}H_{B}}$ to the initial quantum state $|s\rangle$. After applying these unitaries for $p$ levels (corresponding to $i=1,2,\cdots,p$), the output state of the PQC is
$|\bm{\gamma_{p}},\bm{\beta_{p}}\rangle=\prod_{i=1}^{p}e^{-i\beta_{i}H_{B}}e^{-i\gamma_{i}H_{C}},$ (4)
where $\bm{\gamma_{p}}=(\gamma_{1},...,\gamma_{p})$ and $\bm{\beta_{p}}=(\beta_{1},...,\beta_{p})$ are $2p$ variational QAOA parameters. The expectation value $F(\bm{\gamma_{p}},\bm{\beta_{p}})$ of $H_{C}$ in this variational quantum state is defined as
$F({\bm{\gamma_{p}}},{\bm{\beta_{p}}})=-\langle{\bm{\gamma_{p}}},{\bm{\beta_{p}}}|H_{C}|{\bm{\gamma_{p}}},{\bm{\beta_{p}}}\rangle,$ (5)
which can be calculated by repeated measurements of the output quantum state using the computational basis, and the goal is to search for the global optimal parameters
$(\bm{\gamma_{p}}^{opt},\bm{\beta_{p}}^{opt})=\arg\underset{(\bm{\gamma_{p}},\bm{\beta_{p}})}{\max}F(\bm{\gamma_{p}},\bm{\beta_{p}}).$ (6)
The search for optimal parameters is typically done by starting with some guesses of the initial parameters, where the quantum computer computes the expectation value of the output state and delivers the relevant information to the classical computer. These parameters are optimized by the classical optimizer and then delivered to the quantum computer. The outer parameter optimization is repeated until it meets the termination condition (e.g., the maximal number of iterations or convergence tolerance $\varepsilon$ Ni et al. (2024)). To quantify the quality of the QAOA solution (i.e., measure the difference between the output state of QAOA and the ground state of $H_{C}$), we introduce the approximation ratio (AR)
$r=\frac{F({\bm{\gamma_{p}}},{\bm{\beta_{p}}})}{F_{\text{max}}},$ (7)
where $F_{\text{max}}$ is the negative of the ground state energy of $H_{C}$, with equaling $C_{\text{max}}$. For small graph instances, $F_{\text{max}}$ can be obtained by brute force search. For large graphs, an approximation of $F_{\text{max}}$ can be obtained by executing multiple runs of the greedy algorithm Halldórsson and Radhakrishnan (1994) and selecting the largest independent set found across these runs. AR reflects how close the solution provided by QAOA is to the exact solution, $r\leq 1$, with the value of 1 as the exact solution.
II.3 Review of QAOA+
The original QAOA searches for the ground state of $H_{C}$ from a Hilbert space that includes both feasible and infeasible quantum states. As a result, the output states may not always satisfy the independence constraint and violate one or more constraints Ruan et al. (2023); Finžgar et al. (2023). To overcome this challenge, Hadfield et al. Hadfield et al. (2017, 2019) introduce a novel quantum alternating operator denoted as QAOA+. Different from incorporating the independence constraint into the target Hamiltonian utilizing the Lagrange multiplier, QAOA+ encodes constraints into the mixer Hamiltonian $H_{B}$. This allows QAOA+ to search for the optimal solution within a feasible subspace that only contains feasible quantum states. Consequently, the target Hamiltonian
$H_{C}=\sum_{u=0}^{n-1}\frac{\sigma_{u}^{z}-I}{2}$ (8)
is obtained by converting the $x_{u}$ in $-C_{1}(x)$ to a quantum spin $\frac{I-\sigma_{u}^{z}}{2}$, thereby encoding the MIS solution into the ground state of $H_{C}$. Similar to QAOA, QAOA+ approaches the ground state of $H_{C}$ by alternately applying unitaries $e^{-i\gamma_{i}H_{C}}$ and $e^{-i\beta_{i}H_{B}}$ on an initial quantum state $|s\rangle$ required to be trivial to implement. However, in QAOA+, the initial quantum state must be the superposition of feasible states. This requirement ensures that the feasible subspace contains no states that violate the constraints. In addition, the initial quantum state $|s\rangle$ may not be the ground state of $H_{B}$ in QAOA+, unlike in QAOA.
Figure 2: The implementation of $e^{-i\beta_{i}B_{u}}$ when $d(u)\geq 1$. The multi-qubit controlled RX gate can be decomposed into two multi-qubit controlled gates and a single-qubit controlled RX gate, utilizing an auxiliary qubit Vale et al. (2023). The auxiliary qubit is initialized as $|0\rangle$ and is flipped to $|1\rangle$ only when the quantum states of all adjacent vertices are $|0\rangle$ (i.e., $f(u)=0$). Once this condition is met, an RX gate is applied on the $u$-th qubit. After the operation, the state of the auxiliary qubit can be reset from $|1\rangle$ to $|0\rangle$ using the right multi-qubit controlled gate operation. Otherwise, the auxiliary qubit is always $|0\rangle$ when $f(u)>0$, and there is an Identity operation applied to the $u$-th qubit.
Figure 3: A schematic of directly solving the MIS problem on the target graph using QAOA or QAOA+ ansatz, where the concrete form of the initial quantum state $|s\rangle$, the mixer Hamiltonian $H_{B}$, and the target Hamiltonian $H_{C}$ are relevant to the choose of ansatz.
The design of the mixing Hamiltonian $H_{B}$ must preserve the feasibility of the subspace (i.e., ensure the subspace only has quantum states obeying constraints) and explore a new feasible subspace (i.e., provide transitions between various feasible states) Hadfield et al. (2019). For an independent vertex subset $V_{s}$, deleting an arbitrary vertice $v_{a}$ (i.e., the state of $x_{v_{a}}$ is converted from 1 to 0) from $V_{s}$ does not change its independence. However, a new vertex $v_{a}$ can be added to $V_{s}$ (i.e., flipping $x_{v_{a}}$ from 0 to 1) without destroying its independence only when there are no adjacent vertices of $v_{a}$ in $V_{s}$. Thus, for any node, controlling its bit-flip operation based on the states of its neighbors is sufficient to both add and remove vertices while preserving the independence property Hadfield et al. (2019). The bit-flip of $v_{a}$ occurs only if
$f(v_{a})=\sum_{j=1}^{d(v_{a})}x_{v_{j,a}}=0,$ (9)
where $v_{j,a}$ is the $j$-th adjacent vertex of $v_{a}$, and $d(v_{a})$ is the number of adjacent vertices of $v_{a}$, with $j=1,\cdots,d(v_{a})$. Mapping to a quantum circuit, for any vertex $u$, a Pauli-X operation is applied on the $u$-th qubit (i.e., this qubit is flipped) if and only if all quantum states of adjacent vertices of $u$ are $|0\rangle$ (i.e., $f(u)=0$). Otherwise, a Pauli-I operation is applied on the $u$-th qubit. The corresponding partial mixer Hamiltonian $B_{u}$ can be described as
$\begin{split}B_{u}=\sum_{f(u)=0}|x_{v_{1,u}}\cdots x_{v_{d(u),u}}\rangle\langle x_{v_{1,u}}\cdots x_{v_{d(u),u}}|\otimes\sigma_{u}^{x}\\ +\sum_{f(u)>0}|x_{v_{1,u}}\cdots x_{v_{d(u),u}}\rangle\langle x_{v_{1,u}}\cdots x_{v_{d(u),u}}|\otimes I.\end{split}$ (10)
The mixer Hamiltonian is given by $H_{B}=\sum_{u=0}^{n-1}B_{u}$. Specifically, the quantum state resulting from applying $e^{-i\beta_{i}B_{u}}$ to the single feasible quantum state $|x_{o}\rangle$ is
$e^{-i\beta_{i}B_{u}}|x_{o}\rangle=\begin{cases}&RX(2\beta_{i})|x_{o}\rangle,f(u)=0\\ &|x_{o}\rangle,f(u)>0,\end{cases}$ (11)
where the detailed result of $RX(2\beta_{i})|x_{o}\rangle$ is $\cos 2\beta_{i}|x_{o}\rangle-i\sin 2\beta_{i}|x_{0}\cdots\bar{x_{u}}\cdots x_{n-1}\rangle$, and $\bar{x_{u}}$ denotes the state of $u$-th qubit is converted from $|0\rangle$ to $|1\rangle$ (or ($|1\rangle$ to $|0\rangle$), implementing the addition or removal of node $u$ without generating infeasible quantum states. The quantum circuit for $e^{-i\beta_{i}B_{u}}$, shown in FIG. 2, corresponds to a multi-qubit controlled RX gate when the node has $d(u)\geq 1$ adjacent vertices, and to a single RX gate when $d(u)=0$. To help readers understand our subsequent work better, FIG. 3 provides a schematic illustrating the process of directly solving the MIS problem on the target graph using QAOA or QAOA+. Additionally, Table 1 is provided to clearly illustrate the differences between using QAOA and QAOA+ to solve the MIS problem.
Table 1: The difference between QAOA and QAOA+ when solving the MIS problem
QAOA QAOA+
Manipulation of constraint Add it into $H_{C}$ by Lagrange multiplier Encode it into $H_{B}$
Initial state $|s\rangle$ $|s\rangle=|+\rangle^{\otimes n}$ The superposition of feasible solutions
Mixer Hamiltonian $H_{B}$ $H_{B}=-\sum_{u=0}^{n-1}\sigma_{u}^{x}$ $H_{B}=\sum_{u=0}^{n-1}B_{u}$
Target Hamiltonian $H_{C}$ $H_{C}=\sum_{u=0}^{n-1}\frac{\sigma_{u}^{z}-I}{2}+\lambda\sum_{(u,v)\in E}\frac{I-\sigma_{u}^{z}-\sigma_{v}^{z}+\sigma_{u}^{z}\sigma_{v}^{z}}{4}$ $H_{C}=\sum_{u=0}^{n-1}\frac{\sigma_{u}^{z}-I}{2}$
$|s\rangle$ is the ground state of $H_{B}$ Yes No requirement
Hilbert space Including feasible and infeasible states Feasible states
III PROGRESSIVE QUANTUM ALGORITHM
Any arbitrary graph has at least one induced subgraph with the same MIS solution as it. Based on this property, PQA iteratively constructs a derived subgraph that includes the MIS solution of $G$ but is smaller in graph size. Naturally, solving the MIS problem on this subgraph requires fewer resources than directly solving it on the original graph. However, multiple derived subgraphs may exist, each with a different size but sharing the same MIS solution as $G$. As a result, various numbers of quantum gates and qubits are needed to implement the $p$-level QAOA+ ansatz for each subgraph. Fewer quantum gates and qubits are required in the realization of $e^{-i\beta_{i}H_{B}}$ and $e^{-i\gamma_{i}H_{C}}$ as the number of nodes and edges decreases. Especially, in the realization of $e^{-i\beta_{i}H_{B}}$, multiple $d(v)$-qubit controlled RX gates may be required. The overhead of implementing the $d(v)$-qubit controlled RX gate is costly but decreases when the node $v$ has fewer neighborhood nodes (i.e., a smaller $d(v)$) Vale et al. (2023); Tomesh et al. (2022b). To minimize the required qubits and reduce the overhead of implementing multi-qubit controlled RX gates, an ideal subgraph should not only have the same MIS solution as the target graph but also be small and sparse (i.e., with fewer edges and where the number of edges is less than the square of the number of nodes). Constructing an ideal subgraph is a key aspect of PQA. The following sections introduce the rules for constructing induced subgraphs and methods for estimating whether they share the same MIS solution as the target graph.
III.1 The construction rules of the subgraph
To construct such an ideal subgraph, PQA starts from a sparse and small-sized initial subgraph $G_{0}$ and gradually expands it. During each expansion, the process favors adding the node with the lowest “closeness” to minimize the effect of the new node on the sparsity of the subgraph. The closeness is measured by a classical objective function
$C(v_{c})=\sum_{v_{c}\in V_{c},v_{q}\in V_{q}}E(v_{c},v_{q}),$ (12)
where $V_{c}=V-V_{q}$ is the vertex candidate set, and $V_{q}$ is the vertex set of the current induced subgraph $G_{q}$. If the nodes $v_{c}$ and $v_{q}$ are adjacent (i.e., $(v_{c},v_{q})\in E$), $E(v_{c},v_{q})=1$. Otherwise, $E(v_{c},v_{q})=0$. The value of $C(v_{c})$ measures how the addition of vertex $v_{c}$ to the current subgraph $G_{q}$ affects its sparsity, and a small value of $C(v_{c})$ indicates a minimal change in sparsity. The choice of $G_{0}$ is not random, and it also involves the selection of the initial node in addition to graph expansion (i.e., the addition of new vertices and edges). The detailed selection rules are as follows.
Figure 4: A procedure schematic for PQA. PQA starts with an initial subgraph $G_{0}$ and solves the MIS problem using $p$-level QAOA+ ansatz, where the parameters of PQC are randomly initialized. Then, PQA progressively enlarges the graph and solves the MIS problem on the updated subgraph, retraining the PQC with pre-trained parameters. Notably, PQA incorporates two early exit mechanisms to terminate ineffective optimization, conserving quantum resources. The first mechanism assesses whether the current optimization process should proceed. The second mechanism evaluates whether further graph expansion is necessary. If the first mechanism is triggered, the PQC outputs the quantum state obtained from the previous subgraph $G_{i-1}$. The output state that corresponds to the maximal $F_{G_{j},p}$ ($j=0,1,...,i$) is given when the second mechanism is triggered. Otherwise, the quantum state is obtained from the current subgraph $G_{i}$ when the stop condition is met.
The choice of the first node. To construct a sparse initial subgraph $G_{0}$, the node with a minimum degree in the target graph $G$ is chosen as the first node of $G_{0}$. If there are $l(>1)$ nodes with a minimum degree, one is randomly chosen as the first node. The extension of subgraph. The addition of new nodes should have minimal impact on graph sparsity. That is, the node that has $\min C(v_{c})$ in the candidate set is the next vertex to join the current subgraph. If there are $l(>1)$ nodes with $\min C(v_{c})$, further calculations are performed. Specifically, for each node with $\min C(v_{c})$, we assume it is added to the current subgraph and calculate its corresponding $\min C(v_{c})$, resulting in $l$ calculation results. The node with the minimum value among these results is chosen to expand the current subgraph. If multiple nodes meet this criteria, one is randomly selected as the next node. During the extension, new edges may be introduced. Specifically, any edge $(v_{q},v_{new})$ should be added to the current subgraph if the newly added node $v_{new}$ is adjacent to existing nodes $v_{q}$, resulting in an expanded subgraph. Additionally, an example is provided in Appendix A to help readers understand the process of graph expansion.
III.2 Progressive quantum optimization
After constructing an initial induced subgraph $G_{0}$, PQA builds a parameterized quantum circuit using the $p$-level QAOA+ ansatz to solve the MIS problem on $G_{0}$, with the parameters randomly initialized. After optimization, PQA continues to expand the graph size according to the given expansion rules and constructs $p$-level QAOA+ ansatz for the new subgraph to solve the MIS problem. At this stage, the PQC reuses the optimized QAOA parameters from the previous subgraph and re-optimizes these pre-trained parameters. This process is referred to as “parameter transfer” Galda et al. (2021); Shaydulin et al. (2023); Montanez-Barrera et al. (2024); Wurtz and Lykov (2021); Brandao et al. (2018); Li et al. (2023). After each optimization round, PQA gets an expectation function value $F_{G_{i},p}$ about the subgraph $G_{i}$, and PQA terminates if the values satisfy $\Delta_{1}\leq\xi$ and $\Delta_{2}\leq\xi$ or if the size matches that of the target graph in the worst case, where
$\Delta_{1}=\lvert F_{G_{i+1},p}-F_{G_{i},p}\rvert,$ (13)
$\Delta_{2}=\lvert F_{G_{i},p}-F_{G_{i-1},p}\rvert.$
Here, $\xi$ is the tolerance for function values, and $F_{G_{i},p}$ ($F_{G_{i-1},p}$) represents the function value obtained using $p$-level QAOA+ ansatz to solve the MIS problem on the current (previous) subgraph $G_{i}$ ($G_{i-1}$). Here, we explain why $\Delta_{1}\leq\xi$ and $\Delta_{2}\leq\xi$ are used as the termination conditions. According to the properties of the MIS problem, adding a new vertex to the derived subgraph $G_{i}$ can cause the newly constructed subgraph $G_{i+1}$ to have an identical or larger independence number compared with $G_{i}$. In other words, $\beta(G_{i})\leq\beta(G_{i+1})$ (i.e., $F_{\text{max},G_{i}}\leq F_{\text{max},G_{i+1}}$). When PQA starts with a subgraph that has the same MIS solution as the target graph, the growth in the maximal expected function values $F_{\text{max},G_{i}}$ will cease as the subgraph grows (i.e., $F_{\text{max},G_{i}}$ reaches a steady value). Nevertheless, the value of $F_{\text{max},G_{i}}$ might also ascend with the increase of the graph size after a brief period of stagnation when PQA starts with a different initial subgraph. During the transient stationary phase, the addition of new nodes provides no benefit (i.e., without incrementing the maximal expectation function value). This occurs due to being trapped in a local subgraph, which is affected by sub-optimal node selection. For the latter case, the subsequent enlargement of the graph might enhance performance, but it concurrently requires more quantum resources to construct the corresponding quantum circuits. At this time, scaling the graph size without regard to resources is not advisable, and PQA should terminate the process in a timely manner. Given these considerations, we prioritize the stable phase of change in the expected function value during optimization, and PQA is halted if the expected function value remains stable over three consecutive optimization rounds. By this stage, we may have discovered a derived subgraph that includes the MIS solution of the target graph, obtaining an approximate or exact solution by solving the MIS problem on this subgraph. The quality of reused parameters can affect the optimization result for each subgraph. During actual optimization, the expectation function value may oscillate back and forth with increasing subgraph size when the parameter transfer performs poorly, delaying the fulfillment of the stop condition, such as shown in Appendix B. In such cases, PQA may not satisfy the termination condition and continue to execute optimization until reaching the same scale as the target graph, even though a subgraph with the same solution as the target graph has been constructed. This contradicts the original goal of PQA. To reduce running costs in PQA, we first aim to improve the quality of reused parameters by a greedy approach, as done in Ref. Shaydulin et al. (2019); Sack et al. (2023), when solving the MIS problem of $G_{0}$. Specifically, we can execute $c$ rounds of optimization when PQA solves the MIS problem on the initial induced subgraph, then select one set of the optimized parameters with the maximal function value in $c$ sets of results. The preserved parameters are reused to solve the MIS problem in the next subgraph. In addition, early exit mechanisms are required to end the low-cost optimization when appropriate. These mechanisms are based on evaluating the quality of the reused parameters and the optimization results for each subgraph. After each graph extension, we first decide whether to proceed with the optimization round by assessing the quality of the reused (initial) parameters. Generally, PQA can converge to quasi-optima after a few iterations (or even no optimization) when it starts with high-quality initial parameters. Otherwise, it may require abundant iterations to converge or get stuck in a low-quality local solution. Thus, we terminate PQA early and skip the current optimization round if the expected function value, based on the initial parameters, is significantly lower than that obtained from the previous subgraph’s optimization. If the quality of the initial parameters meets the rules we set, a new round of early exit judgments is carried out after the optimization. If the latest expected function value is significantly lower than the maximal value obtained from previous subgraphs, it indicates that the optimization has fallen into a poor local solution. At this time, PQA will halt further graph expansion and optimization, returning the maximal function value in this run and the corresponding output state. Otherwise, PQA continues to execute if it satisfies the above settings. In summary, to reduce the resources required during a PQA run, we introduced two early exit mechanisms that determine whether subsequent optimization rounds are necessary based on ‘flags’ (the quality of the initial parameters and the optimization results). Notably, PQA cannot get the optimal solution to the problem through an optimization run, and it simply aims to obtain a quasi-optimal solution with as few quantum resources as possible. More detailed procedures of PQA are shown in FIG. 4. PQA is a hybrid quantum-classical algorithm, where this concept encompasses two aspects. Firstly, when the QAOA+ algorithm is used to solve the MIS problem on a subgraph, the classical computer handles parameter optimization, while the quantum computer computes the expected function value. This constitutes an external parameter optimization loop, achieving synergy between classical and quantum computation. Secondly, in constructing the subgraph, the classical computer creates a new subgraph based on the structural information of the target graph. Meanwhile, the quantum computer solves the MIS problem on the subgraph and informs the classical computer whether to continue expanding the subgraph based on the optimization results. This forms another collaborative process. In summary, PQA leverages the parallel computing capabilities of the quantum computer to solve the MIS problem on each subgraph, while sharing part of the costs (such as obtaining target graph information and parameter adjustment) with the classical computer, rather than having either the quantum or classical computer handle all resource overheads alone.
IV RESULTS
In our work, the effects of graph construction rules and the initial graph size $n_{0}$ on the performance of PQA are first investigated in Section IV.1. Then, we compare the performance of different strategies on various graphs, with more details in Section IV.2.
IV.1 The effects of graph construction rules and $n_{0}$ on the performance of PQA
To investigate the impact of graph construction rules and the initial graph size $n_{0}$ on the performance of PQA, we randomly generate multiple ER graphs. For the given graph instances, we conduct $100\times p$ runs of PQA following various graph construction rules, where $n_{0}=1$, $\xi=0.1$, and the quantum state $|s\rangle$ is initialized as $|\mathbb{0}\rangle$ at each solving process. Subsequently, we calculate the achieved AR by PQA at level $p=1$ under different construction rules. Notably, the AR is the ratio between the expectation function value obtained by PQA on the final induced subgraph and $F_{\text{max},G}$. Here, $F_{\text{max},G}$ denotes the maximal expectation function value on the target graph, equivalent to the independence number of the target graph. The AR is less than or equal to one, and an AR of one signifies that PQA achieves the MIS solution of the target graph by solving the MIS problem on the current subgraph.
IV.1.1 The effects of construction rules on the performance of PQA.
For each graph instance, we first calculate the optimal and average approximation ratio (OAR, AAR) obtained by multiple optimization runs of PQA when it constructs subgraphs according to different rules. Then, we average the obtained OAR and AAR across all graph instances to get the overall average OAR and AAR for PQA. From these simulation results given in FIG. 5, we observe that the expansion rules play a more significant role in the performance of PQA compared with the choice of the first node. Specifically, PQA achieves a much higher AR when following our expansion rules instead of random expansion. Given these phenomena, we analyze the variation of the expected function value during optimization and find that the parameter transfer tends to perform poorly (i.e., getting stuck in a low-quality local solution) when random vertex additions cause significant changes in the graph structure, leading to early termination of the optimization run. This may occur because significant changes in the graph structure lead to drastic shifts in the function landscape, rendering the transferred parameters suboptimal for the new landscape. In addition, FIG. 6 provides the total number of iterations required by PQA under various construction rules while achieving their respective OAR at $p=1$. These findings suggest that PQA requires far fewer iterations to obtain a quasi-optimal solution when following our graph expansion rules, with a $95\%$ reduction in iterations compared with random first node and expansion (i.e., ’NN’). In summary, the results presented in FIG. 5 and FIG. 6 collectively demonstrate the efficiency of our proposed graph construction rules.
Figure 5: The effects of graph construction rules on the AR achieved by PQA, where ‘NN’ suggests PQA arbitrarily chooses the first node and subsequent nodes without any rules. ‘NY’ represents that the first node is randomly chosen, while other newly added nodes obey our predetermined expansion rules.
Figure 6: The effects of graph construction rules on the total consumption of iterations to get the quasi-optimal solution.
Figure 7: The OAR achieved by PQA at various level depths $p$ and $n_{0}$ on given ER graphs, where $n_{0}$ is the number of nodes in the initial induced subgraph $G_{0}$. Each solid dot in the curve represents the AR obtained by PQA at level depth $p$ when it starts from a subgraph with $n_{0}$. Specifically, the curve reflects the AR obtained by DS-QAOA+ under various level depths when $n_{0}=14$.
IV.1.2 The effects of $n_{0}$ on the performance of PQA.
Furthermore, we explore the impact of $n_{0}$ on the performance of PQA under various $p$. In our simulations, we vary the level depth from one to five and perform $100\times p$ runs of PQA at each level. The resulting AR versus the level depth and $n_{0}$ is depicted in FIG. 7. These simulation results suggest that PQA may require more level depths to get the exact solution when the initial problem size is large. Concretely, PQA can achieve the exact MIS solution (i.e., OAR = 1) at $p=1$ when starting with initial subgraphs of size $n_{0}=2,4,6$, while for PQA with $n_{0}=14$ (i.e., DS-QAOA+), its required level depth is at least five. In general, the algorithm often requires multiple optimization runs at specific level depths to obtain the exact solution or reach a given OAR threshold. During parameter optimization, each run consists of numerous iterations, with each iteration requiring multiple calls to the quantum processing units (QPUs) to compute the expectation function value. More iterations increase the demand on QPUs. To comprehensively examine the effect of $n_{0}$ on the performance of PQA, we analyze the total number of optimization runs (TRs) and iterations (TITRs) when PQA initializes from various $n_{0}$ values and achieves a given OAR at the required level depths (RLDs). Specifically, for each graph instance, we first calculate the OAR achieved by multiple runs at each level depth to determine the required level depths for each algorithm to reach the given OAR. Then, we calculate the number of runs $Q$ when the resulting AR exceeds the given OAR threshold over multiple optimization runs. We regard the ratio of $Q$ to the total number of runs $R$ at level depth $p$ as the probability of exceeding that threshold. The reciprocal of this probability gives the TRs required for the algorithm to reach the given OAR. In addition, we calculate the average iterations (AITRs) needed for an optimization run at the RLDs, and the TITRs required for the algorithm to achieve the given OAR at the RLDs equals the product of TRs and AITRs. For each graph instance, we conduct a similar analysis to obtain the RLDs, TRs, and TITRs. For each graph type, we compute the average of multiple RLDs, TRs, and TITRs to estimate the number of level depths, total optimization runs, and total iterations required for the algorithm to achieve a given OAR when solving the MIS problem on a given graph type.
Table 2: The required level depths (RLDs), total number of optimization runs (TRs), total iterations (TITRs), total qubits (T-qubits), total circuit depths (TCDs), total multi-qubit CR (T-MQCR) gates, and total quantum (T-Q) gates for PQA to achieve the given OAR under different values of $n_{0}$.
 OAR $n_{0}$ RLDs TRs TITRs T-qubits TCDs T-MQCR gates T-Q gates
0.9 2 1 1.5562 76.0375 8.7926 23.568 6.6096 17.5852
4 3.2 8.4233 704.2204 74.54 1079.1239 340.6935 716.1204
6 1.6 1.4397 95.0968 11.5176 66.9053 21.2897 43.2461
8 2.4 1.9594 125.2674 20.1942 152.0068 49.0538 98.1076
10 3.4 4.1255 294.5122 49.5827 489.149 158.6944 317.3888
12 4.2 7.1945 549.2766 100.3954 1507.5039 490.7771 981.5542
14 4.8 46.8667 2387.0363 656.13332 14169.9319 4613.4662 9226.9324
0.85 2 1 1.5562 76.0375 8.7926 23.568 6.6096 17.5852
4 3 9.7833 832.7083 94.7 1252.4839 397.3335 829.4004
6 1.4 1.5901 98.982 12.7205 67.6558 21.5299 43.7264
8 2.2 1.9921 125.2955 20.8618 150.6104 48.6473 97.2946
10 3.2 4.783 371.4308 60.6338 584.0127 189.7596 379.5192
12 4.0 5.5698 426.9543 77.7713 1125.6186 366.4565 732.913
14 4.6 33.0571 1521.2784 462.8 7244.8851 2358.7998 4717.5996
Moreover, for each graph instance, the total number of qubits (T-qubits), total circuit depths (TCDs), total multi-qubit controlled RX (T-MQCR) gates, and total quantum (T-Q) gates demanded by each algorithm are also calculated when they achieve the given OAR. Combining with these indicators, we aim to clearly reflect the resource expenses for PQA with different $n_{0}$. The circuit depth measures how many “layers” of quantum gates are executed in parallel, and all quantum gates that can be applied simultaneously in the circuit are regarded as one layer. The circuit depth reflects the execution time of the quantum calculation process Ni et al. (2024); Guerreschi and Matsuura (2019). Deeper circuits require more time for execution, increasing the risk of noise and errors Du et al. (2022). In this work, the circuit depth is obtained by directly using the DAGCircuit interface in MindSpore Quantum Developer (2021). The gate resource refers to the number and types of quantum gates required in a quantum circuit, such as single-qubit rotation gates and multi-qubit controlled rotation (CR) gates, with the latter being more complex to implement on quantum hardware. Gate resources reflect the operational complexity of quantum circuits, with more gates generally leading to greater complexity. In our work, we count the number of single-qubit RX gates, single-qubit RZ gates, and multi-qubit controlled RX gates. The sum of these gives the consumption of quantum gates required by the PQC with the $p$-level QAOA+ ansatz. In our calculations, the number of $d(v)$-qubit controlled RX gates in one level QAOA+ ansatz is incremented by one if $d(v)\geq 1$. The total quantum resources are the product of the average consumption per run and TRs at their respective RLDs. From these simulation results shown in Table 2, we observe that PQA’s resource demands increase as it starts with a larger initial induced subgraph. Specifically, PQA with $n_{0}=2$ achieves the given OAR while consuming only about $4.718\%$ of the quantum resources required by DS-QAOA+ (i.e., PQA with $n_{0}=14$). This represents a savings of approximately $95.9\%$ of iterations compared with DS-QAOA+, revealing substantial efficiency gains with PQA. In conclusion, these above investigations emphasize the sensitivity of PQA to the choice of $n_{0}$. Its performance may potentially decline when initialized with a large-scale initial subgraph. To maintain robust performance in subsequent simulations, we opt for $n_{0}=2$. This decision is informed by our observation that smaller values of $n_{0}$ tend to yield more reliable results, ensuring the stability and effectiveness of PQA across varied scenarios.
(a) OAR, ER graphs with prob = 0.4
(b) OAR, ER graphs with prob = 0.5
(c) OAR, 2-regular graphs
(d) OAR, 3-regular graphs
Figure 8: The AR achieved by PQA, QLS, and DS-QAOA+ under the same level depth and optimization runs. In each subplot, the curves depict the AR achieved by the different algorithms across various graph types as the level depth $p$ increases.
IV.2 The performance comparison among different algorithms
IV.2.1 The OAR obtained by different algorithms under the same number of optimization runs and level depth.
In our simulations, we randomly generate 20 graphs each for different ER graphs (with probability = 0.4, 0.5) and $k$-regular graphs ($k=2,3$) with the number of nodes $n=14$. Subsequently, we respectively test the performance of PQA, QLS, and DS-QAOA+ on these graphs. For PQA, the tolerance on function values $\xi=0.1$ and the quantum state $|s\rangle$ is initialized as $|\mathbb{0}\rangle$ at each solving process. The classical outer parameter optimization loop ends when the convergence tolerance Ni et al. (2024) satisfies $\varepsilon\leq 0.001$. We set the level depth from one to five, and each algorithm is executed $100\times p$ optimization runs at each depth. At the end of the runs, we calculate the OAR achieved by each algorithm. These experimental procedures yield 20 sets of OAR for each graph type. The mean value of these sets provides the average OAR for various algorithms at depth $p$. From these numerical results illustrated in FIG. 8, we observe that the achieved OAR tends to increase as the level depth increases, thereby enhancing the quality of the obtained solution. More importantly, these numerical results demonstrate that the OAR achieved by PQA is superior to that obtained by DS-QAOA+ and QLS at the same level depth and number of optimization runs. More specifically, on the ER graphs with prob = 0.4, the OAR achieved by PQA is approximately $8.24\%$ ($13.668\%$) higher than that achieved by QLS (DS-QAOA+). On the ER graphs with prob = 0.5, the OAR achieved by PQA is approximately $8.128\%$ ($14.422\%$) higher than that achieved by QLS (DS-QAOA+). On the 2-regular graphs, the OAR achieved by PQA is about $3.612\%$ ($4.514\%$) higher than that achieved by QLS (DS-QAOA+). On the 3-regular graphs, the OAR obtained by PQA is about $8.486\%$ and $9.438\%$ higher than that achieved by QLS and DS-QAOA+, respectively.
Table 3: The required level depths (RLDs), the total number of optimization runs (TRs) and iterations (TITRs), total qubits (T-qubits), total circuit depths (TCDs), total multi-qubit CR (T-MQCR) gates, and total quantum (T-Q) gates consumed by different algorithms on specific graph types when they achieve OAR of 0.95.
Types Algo. RLDs TRs TITRs T-qubits TCDs T-MQCR gates T-Q gates
ER, prob = 0.4 DS-QAOA+ 4.9 91.6192 5664.43 1282.669 19304.17 6285.077 12570.15
QLS 4.4 26.8412 1666 371.46 5282.35 1648.8952 3297.79
PQA 1.35 17.7572 1652 112.76 464.35 135.84 337.3868
ER, prob = 0.5 DS-QAOA+ 6.85 339.1095 21624 4747.533 99884.7 32520.6 65041.2
QLS 4.55 34.3853 1939.38 478.59 7011.16 2186.785 4373.57
PQA 1.6 15.8492 1241.81 110.94 527.78 160.869 361.362
2-regular DS-QAOA+ 3.35 10.4152 735.41 145.813 1500.31 488.473 976.946
QLS 2.35 5.35 374.34 50.02 655.375 202.498 405
PQA 1 5.15 336.96 41.74 91.89 22.497 83.48
3-regular DS-QAOA+ 6.3 493.9459 36313 6915.243 133809.9 43566.03 87132.06
QLS 6.21 50.3033 3519.34 690.37 14005.5 4359.867 8719.735
PQA 1.35 20.5773 2319.783 150.21 482.54 125.522 407.431
Table 4: The ratio of quantum resources saved by PQA relative to DS-QAOA+ and QLS when an OAR of 0.95 is achieved.
Types Algo. RLDs TRs TITRs T-qubits TCDs T-MQCR gates T-Q gates
ER, prob = 0.4 DS-QAOA+ 0.7245 0.8062 0.7083 0.9121 0.9759 0.9784 0.9732
QLS 0.6932 0.3384 0.0084 0.6964 0.9121 0.9176 0.8977
ER, prob = 0.5 DS-QAOA+ 0.7664 0.9533 0.943 0.9766 0.9947 0.9951 0.9944
QLS 0.6484 0.5391 0.36 0.7682 0.9247 0.9264 0.9174
2-regular DS-QAOA+ 0.7015 0.5054 0.5418 0.7137 0.9388 0.9539 0.9145
QLS 0.5745 0.0379 0.1 0.1655 0.8598 0.8889 0.7939
3-regular DS-QAOA+ 0.7857 0.9583 0.9361 0.9783 0.9964 0.9971 0.9953
QLS 0.7826 0.5909 0.3408 0.7824 0.9655 0.9712 0.9533
IV.2.2 The quantum resources consumed by various algorithms when they achieve the same OAR.
To provide a more comprehensive comparison of the performance of DS-QAOA+, QLS, and PQA, we further investigate the total number of optimization runs and iterations required by each algorithm to reach a given OAR at the required level depths. Moreover, for each graph instance, the total number of qubits, total circuit depths, and other resources demanded by each algorithm are also calculated. These indicators are used to clearly demonstrate the resource consumption of each algorithm when they reach the same OAR. In the numerical simulation experiments, we set the level depth from one to $n$, and the OAR obtained by each algorithm under multiple runs to be at least $0.95$. We perform 100 optimization runs at each level depth $p$ when $p\leq 5$, and $2^{p}$ optimization runs when $p>5$. For each graph instance, we calculate the RLDs, TRs, ITRs, and other resources when they achieve the OAR of 0.95. Averaging the data of the same graph type provides an estimate of the resources consumed by different algorithms when solving the MIS problem on various graph types to achieve the given OAR. The specific numerical results are shown in Table 3. These results suggest that PQA is usually able to achieve the OAR of 0.95 with fewer resources. Table 4 summarizes the quantum resources saved by PQA relative to DS-QAOA+ and QLS when an OAR of 0.95 is reached on a given graph type. For example, when solving the MIS problem on a 2-regular graph, the total qubits consumed by PQA is only $28.63\%$ of DS-QAOA+, saving $71.37\%$ in qubit usage. These results demonstrate that PQA not only achieves the given OAR of 0.95 in fewer optimization runs and iterations but also significantly reduces the consumption of qubits, circuit depth, and quantum gates compared with DS-QAOA+ and QLS, highlighting its efficiency. By minimizing the quantum resources required while maintaining high performance, PQA may be increasingly promising for practical implementation in scenarios where resource availability is limited. In our work, we also test other OAR thresholds (i.e., 0.8, 0.85, and 0.9) and show the corresponding resources required by each algorithm in the Appendix C.
(a) given OAR = 0.8
(b) given OAR = 0.85
(c) given OAR = 0.9
Figure 9: The required running time of DS-QAOA+, QLS, and PQA when they achieve the same given OAR.
IV.2.3 The comparison of the running time when DS-QAOA+, QLS, and PQA achieve the same OAR.
To assess the performance of different algorithms in achieving the given OAR, we examine the resource consumption of QLS, PQA, and DS-QAOA+. Experimental results show that QLS and PQA typically require fewer resources than DS-QAOA+ to achieve the same OAR. For QLS and PQA, classical computations (i.e., determining whether an edge exists between two points) are involved in updating the global solution and graph expansion. In contrast, DS-QAOA+ excludes such computations. Although numerical simulations show that the time for a single edge judgment is minimal, we include this classical computation time in our runtime comparison for fairness. For QLS and PQA, the time required for a single optimization run includes both quantum and classical components. The classical time $T_{c}$ is the time consumed in determining whether there is an edge between vertices during updating the global solution or expanding the subgraph. Assuming that the time required for a single determination is $t_{j}$, and the total number of determinations required by the algorithm in one run is $TJ$, then
$T_{c}=TJ\times t_{j}.$ (14)
The quantum running time $T_{q}$ includes the time required for initial state preparation, quantum state measurements, and quantum gate evolution across multiple iterations Guerreschi and Matsuura (2019). In each iteration, $M$ repeated measurements are needed to calculate the average expectation value of the objective function. The time consumed in a single repetition is
$t_{q}=t_{p}+t_{M}+(\text{circuit depth})\times t_{G},$ (15)
where $t_{p}$ is the initial state preparation time, $t_{M}$ is the time for a single measurement, $t_{G}$ is the time for a single quantum gate evolution, and circuit depth is the depth of the parameterized quantum circuit. The time consumed for a single iteration is
$T_{ITR}=t_{q}\times M,$ (16)
and the quantum time consumed in a single optimization run is
$T_{q}=T_{ITR}\times\text{ITRs},$ (17)
where ITRs is the number of iterations required for the objective function to converge in this run. Therefore, the time consumed in a single run for QLS and PQA is
$T_{PQA}=T_{q,PQA}+T_{c,PQA},$ (18)
$T_{QLS}=T_{q,QLS}+T_{c,QLS},$ (19)
respectively, while for DS-QAOA+, its running time is
$T_{DS}=T_{q,DS}.$ (20)
This study explores the runtime required by different algorithms to achieve the same OAR across various graph types. The calculations employ the same realistic assumptions as Ref. Guerreschi and Matsuura (2019), namely, $t_{p}+t_{M}$ is one millisecond, and $t_{G}$ is 10 nanoseconds. The tests are conducted on a server equipped with an Intel(R) Xeon(R) Sliver 4316 CPU running at 2.3GHz and 251GB of RAM, where $t_{j}$ is determined to be 0.31226 milliseconds. Numerical simulations in FIG. 9 show that when the given OAR is 0.8 (0.85, 0.9), the runtime of PQA on certain graph types may be slightly higher than that of QLS and DS-QAOA+. However, as shown in Table A1 in Appendix C, PQA requires only a fraction of the qubits to achieve the given OAR compared with QLS and DS-QAOA+. Moreover, as the given OAR increases (i.e., as the quasi-optimal solutions approach the true solution), PQA consumes less runtime than DS-QAOA+ and QLS, while also requiring fewer qubits, multi-qubit controlled rotation gates, and other resources. Numerical results suggest that PQA exhibits various advantages at different OARs and graph types, such as reduced qubit consumption, shorter runtime, or both. In summary, while PQA may not always be the optimal choice for all OARs and graph types, it likely achieves better results with less runtime and resource consumption compared to DS-QAOA+ and QLS when targeting higher-quality solutions.
V CONCLUSIONS
This paper proposes PQA to reduce quantum resource consumption in the application of QAOA+ for solving the MIS problem. The core idea of PQA is to progressively construct a subgraph containing the MIS solution. As a result, solving the MIS problem on this subgraph provides the MIS solution for the target graph. By doing so, PQA can achieve a quasi-optimal solution with reduced resource consumption compared with conventional approaches like DS-QAOA+. Thus, leveraging PQA on NISQ devices may enable solving larger and more complex optimization problems within current resource constraints. Our analysis examines the impact of $n_{0}$ (the number of vertices in the initial subgraph) on both the final AR achieved by PQA and its quantum resource consumption. The results suggest that starting with a large initial subgraph may lead to a performance decline in PQA. To simplify, PQA starts with a small subgraph and adds one node during each graph extension. However, this incremental approach may introduce unnecessary optimization rounds, increasing the iteration count. To overcome this limitation and streamline the optimization process, future research should explore more effective graph construction methods and the optimal choice of initial size $n_{0}$, taking into account the graph structure to maintain PQA performance. A potential way to reduce optimization rounds is by adding multiple vertices during each expansion, but further investigation is needed to understand its impact on PQA performance. Another promising approach is integrating machine learning to optimize the selection of $n_{0}$ and dynamically adjust the graph expansion strategy based on real-time feedback. These advancements could significantly reduce iterations in the external optimization loop, improving the overall efficiency and scalability of PQA for quantum optimization problems. The graph structure plays a role in determining the performance of PQA. Designing more refined and efficient graph expansion rules for specific graph types is essential. This could further enhance the performance of PQA in future applications. Notably, PQA has potential applications in a broader range of constrained combinatorial optimization problems, although this paper focuses on the MIS problem as an example. When extending PQA to tackle various optimization problems, it is crucial to investigate the optimal subgraph construction scale for specific problems. Thus, future research should focus on exploring optimal subgraph construction rules for different problem instances. These investigations may provide valuable insights for fine-tuning PQA and improving its performance and applicability across various combinatorial optimization problems.
ACKNOWLEGEMENTS
This work is supported by National Natural Science Foundation of China (Grant Nos. 62371069, 62372048, 62272056), Beijing Natural Science Foundation (Grant No. 4222031) and BUPT Excellent Ph.D. Students Foundation(CX2023123).
Appendix A The extension of the subgraph
In this appendix, we introduce the process of the graph extension based on two different derived subgraphs as shown in FIG. A1.
Figure A1: The target graph and its induced subgraphs.
Denote the current subgraph as $G_{i}=(V_{i},E_{i})$. For the first subgraph in FIG. A1, $V_{i}=\left\{0,1,2\right\}$ and $E_{i}=\left\{(0,1),(0,2)\right\}$. The candidate vertex set $V_{C}=V-V_{i}=\left\{3,4\right\}$. Our goal is to select the vertex from $V_{C}$ that has the least connection with the current subgraph to be added to $G_{i}$. Specifically, if node 3 is added to $G_{i}$, $C(v_{c}=3)=1$. Similarly, if node 4 is added, $C(v_{c}=4)=1$. Since both nodes have the same connection, we proceed to the next round of selection. When node 3 is added to $V_{i}$, the edge $(0,3)$ is included in $E_{i}$, leaving $V_{C}=\left\{4\right\}$ and $C(v_{c}=4)=2$. Similarly, if node 4 were added, $C(v_{c}=3)=2$. After evaluating both nodes, since they equally satisfy the selection criteria, one can be randomly added to $V_{i}$. If node 3 (or 4) is added, the edge $(0,3)$ (or $(0,4)$) is included in $E_{i}$. The resulting subgraph becomes is $G_{i+1}=\left\{\left\{0,1,2,3\right\},\left\{(0,1),(0,2),(0,3)\right\}\right\}$ (or $G_{i+1}=\left\{\left\{0,1,2,4\right\},\left\{(0,1),(0,2),(0,4)\right\}\right\}$). For the second subgraph in FIG. A1, $V_{i}=\left\{0,1,3\right\}$ and $E_{i}=\left\{(0,1),(0,3)\right\}$. The candidate vertex set $V_{C}=\left\{2,4\right\}$, with $C(v_{c}=2)=1$ and $C(v_{c}=4)=2$. We choose the node with the smallest $C(v_{c})$ value from the candidate vertex set. Thus, node 2 is the next node to be added to $V_{i}$, and the edge $(0,2)$ is added to $E_{i}$. The newly constructed subgraph is $G_{i+1}=\left\{\left\{0,1,2,3\right\},\left\{(0,1),(0,2),(0,3)\right\}\right\}$.
Appendix B The variation of expectation function values in optimization runs of PQA
Figure A2: The variation of expectation function values in optimization runs of PQA at level depth $p$, where $n_{0}=2$ and the maximal expectation function value is 4. In each subplot, each curve corresponds to an optimization run of PQA, where the square on each curve is the obtained function value after each optimization round. The X-axis represents the number of nodes in the current induced subgraph, where the maximal value of the X-axsis is the maximal size of the final subgraph when PQA terminates. (a) $p=4$, (b) $p=5$.
In our simulations, we use PQA (without early stop mechanisms) to solve the MIS problem on an ER graph with $n=14$ vertices, setting the level depths to $p=4$ and $p=5$. At each level depth, PQA performs $20\times p$ optimization runs. This section provides detailed analysis of the changes in the expectation function value during each optimization run of PQA. From the simulation results in FIG. A2, we observe that fluctuations in the expectation function values can delay the termination of PQA. That is, more optimization rounds are required in an optimization run when the expectation function values obviously fluctuate back and forth (i.e., the values do not meet the stop condition). As a result, the final induced subgraph may reach the same size as the target graph when the optimization process of PQA concludes.
Table A1: The required level depths (RLDs), the total number of optimization runs (TRs) and iterations (TITRs), and other quantum resources consumed by different algorithms on specific graph types when they achieve various given OAR.
 OAR Types Algo. RLDs TRs TITRs T-qubits TCDs T-MQCR gates T-Q gates
0.8 ER, prob = 0.4 DS-QAOA+ 2.75 7.196 333.8991 100.74 850.93 277.05 554.09
QLS 2.4 8.6584 411.3193 119.82 916.92 289.46 578.93
PQA 1.15 12.2964 869.1626 76.24 271.75 81.16 190.59
ER, prob = 0.5 DS-QAOA+ 2.85 37.5934 1745.184 526.31 7678.45 2499.96 4999.92
QLS 2.2 15.0766 619.6376 209.81 1466.2 462.84 925.67
PQA 1.25 6.385 586.3698 45.65 170.48 51.4 119.4
2-regular DS-QAOA+ 1.05 1.213 59.7134 16.98 54.77 17.83 35.66
QLS 1.1 1.2337 88.2352 11.54 43.55 13.26 26.52
PQA 1 1.0415 67.185 8.44 18.58 4.54 16.88
3-regular DS-QAOA+ 1.2 7.2102 316.0523 100.94 372.05 121.13 242.26
QLS 1.1579 8.0593 338.0899 110.61 412.3 128.42 256.85
PQA 1 1.0683 62.5453 7.69 17.95 4.59 15.38
0.85 ER, prob = 0.4 DS-QAOA+ 3.75 18.2939 928.8368 4674.59 2949.89 960.43 1920.86
QLS 3.15 14.6076 798.6345 226.75 2037.03 641.73 1283.47
PQA 1.2 21.0906 1782.131 104.04 523.05 158.18 362.76
ER, prob = 0.5 DS-QAOA+ 4.85 15.6669 854.652 24432.58 3839.96 1250.22 2500.44
QLS 3.35 22.4242 1066.927 282.38 3327.75 1049.37 2098.75
PQA 1.45 15.4684 1108.006 144.04 526.7 161.64 361.96
2-regular DS-QAOA+ 1.15 1.361 74.5434 835.99 67.3 21.91 43.82
QLS 1.25 6.1747 367.6252 48.16 257.79 79.34 158.69
PQA 1 1.1305 72.5766 41.75 20.17 4.94 18.33
3-regular DS-QAOA+ 3.65 15.6257 851.1209 4424.73 2452 798.47 1597
QLS 3.3684 16.6128 913.7158 282.41 2469 778.84 1557.7
PQA 1.3 18.0773 1853.443 150.21 409.45 106.66 345.28
0.9 ER, prob = 0.4 DS-QAOA+ 4.2 39.1584 2228.736 4674.59 60302.18 19633.27 39266.53
QLS 3.6 22.0714 1244.793 226.75 2628.11 823.03 1646.06
PQA 1.35 16.3847 1337.357 104.04 428.46 125.34 311.01
ER, prob = 0.5 DS-QAOA+ 5.8 83.4828 4900.988 24432.58 465266.05 151481.97 302963.94
QLS 3.9 28.0586 1426.722 282.38 3535.17 1105.62 2211.24
PQA 1.5 20.2879 1994.846 144.04 621.82 184.62 444.31
2-regular DS-QAOA+ 2.6 6.8045 389.1411 835.99 6675.96 2173.57 4347.14
QLS 2.35 5.354 374.3412 48.16 459.21 143.97 287.94
PQA 1 5.151 336.9634 41.75 91.9 22.5 83.5
3-regular DS-QAOA+ 4.8 45.5427 2973.713 4424.73 65233.19 21238.71 42477.43
QLS 4.5789 25.5096 1587.143 282.41 4251.92 1328.6 2657.2
PQA 1.3 20.5773 2261.173 150.21 466.08 121.41 393.03
Table A2: The ratio of quantum resources saved by PQA relative to DS-QAOA+ and QLS when a given OAR is reached.
 OAR Types Algo. T-qubits TCDs T-MQCR gates T-Q gates
0.8 ER, prob = 0.4 DS-QAOA+ 0.2433 0.6806 0.7071 0.6560
QLS 0.3638 0.7036 0.7196 0.6708
ER, prob = 0.5 DS-QAOA+ 0.9133 0.9778 0.9794 0.9761
QLS 0.7824 0.8837 0.8889 0.8710
2-regular DS-QAOA+ 0.5029 0.6607 0.7448 0.5266
QLS 0.2682 0.5733 0.6569 0.3635
3-regular DS-QAOA+ 0.9238 0.9518 0.9621 0.9365
QLS 0.9305 0.9565 0.9642 0.9401
0.85 ER, prob = 0.4 DS-QAOA+ 0.4647 0.8227 0.8353 0.8111
QLS 0.3219 0.7432 0.7535 0.7174
ER, prob = 0.5 DS-QAOA+ 0.4746 0.8628 0.8707 0.8552
QLS 0.6308 0.8417 0.8460 0.8275
2-regular DS-QAOA+ 0.5191 0.7003 0.7746 0.5818
QLS 0.8413 0.9218 0.9378 0.8845
3-regular DS-QAOA+ 0.3968 0.8330 0.8664 0.7838
QLS 0.4212 0.8342 0.8631 0.7783
0.9 ER, prob = 0.4 DS-QAOA+ 0.9777 0.9929 0.9936 0.9921
QLS 0.5412 0.8370 0.8477 0.8109
ER, prob = 0.5 DS-QAOA+ 0.9941 0.9987 0.9988 0.9985
QLS 0.4899 0.8241 0.8330 0.7991
2-regular DS-QAOA+ 0.9501 0.9862 0.9896 0.9808
QLS 0.1331 0.7999 0.8437 0.7100
3-regular DS-QAOA+ 0.9661 0.9929 0.9943 0.9907
QLS 0.4681 0.8904 0.9086 0.8521
Appendix C The quantum resources consumed by different algorithms under various given OAR.
We analyze the resource consumption of each algorithm across various graph types and at different OAR thresholds. The focus is on critical factors such as required level depths, the total number of optimization runs, and so on. Numerical results are presented in Table A1. These data show that while PQA typically requires more iterations than QLS to achieve certain OAR thresholds on specific graph types, it consistently consumes fewer qubits, has shallower circuit depths, and uses fewer quantum gates. Table A2 further quantifies these advantages. For example, PQA saves $54.12\%$ of qubits and $81.09\%$ of quantum gates compared to QLS when solving the MIS problem on ER graphs (with a probability of 0.4) and reaching an OAR of 0.9. The numerical simulations reveal that resource consumption varies with the OAR threshold for each algorithm. While PQA does not consistently outperform QLS and DS-QAOA+ across all resource metrics at every threshold, our results show that between OAR thresholds of 0.8 and 0.9, PQA is the preferred choice when a moderate increase in iterations is acceptable in exchange for shallower level depths and fewer qubits. Conversely, if minimizing the number of iterations is the priority, QLS is more efficient for solving the MIS problem. Nonetheless, DS-QAOA+ usually requires more resources than both QLS and PQA.